Bentley Corp. and Rolls Manufacturing are considering a merger. The possible states of the economy and each company’s value in that state are shown here:

   State Probability Bentley Rolls
  Boom .80 $ 310,000 $ 280,000
  Recession .20 $ 120,000 $ 90,000
Bentley currently has a bond issue outstanding with a face value of $135,000. Rolls is an all-equity company.
a.

What is the value of each company before the merger? (Do not round intermediate calculations and round your answers to the nearest whole number, e.g., 32.)

  Value of Bentley $   
  Value of Rolls $   
b.

What are the values of each company’s debt and equity before the merger? (Leave no cells blank - be certain to enter "0" wherever required. Do not round intermediate calculations and round your answers to the nearest whole number, e.g., 32.)

  Equity of Bentley $   
  Debt of Bentley $   
  Equity of Rolls $   
  Debt of Rolls $   
Answer #1

a) Value of Bentley $ 272,000 =310000*80%+120000*20% Value of Rolls $ 242,000 =280000*80%+90000* 20% b) Equity of Rolls $ 242

*Please rate thumbs up